i really dont get this one can u plz help me
Thanks
Solve Problems Using Linear Systems
7.It took a patrol boat 5h to travel 60 km up a river against the current and 3h for the retunr trip with the current. Find the speed of the boat in still water and the speed of the current.

Answer by checkley77(12844)

D=RT OR D/T=R
60/5=12 KM/H IS THE SPEED OF THE BOAT AGAINST THE CURRENT.
60/3=20 KM/H IS THE SPPED OF THE BOAT WITH THE CURRENT.
(20-12)/2=8/2=4 KM/H IS THE SPEED OF THE CURRENT.
12+4=16 KM/H IS THE SPEED OF THE BOAT IN STILL WATER.
20-4=16 KM/H DITTO.